OpenStudy (anonymous):

03 = 3 A. True B. False is it false??

OpenStudy (anonymous):

False.

OpenStudy (anonymous):

I think, .03 would not equal 3

OpenStudy (anonymous):

sorry 0^3

OpenStudy (squirrels):

False

OpenStudy (squirrels):

0^3=0*0*0= 0

OpenStudy (linda3):

O to the power of 3 means 0 x0 x0 which would equal 0

OpenStudy (squirrels):

beatcha to it c;

OpenStudy (linda3):

;-)

OpenStudy (anonymous):

Still false lol

OpenStudy (squirrels):

oh?

OpenStudy (squirrels):

oh yea it is hue ( ͡° ͜ʖ ͡°)